Suppose that the reserve requirement for checking deposits is 20 percent and that banks do not hold any excess reserves.

If the Fed sells $3 million of government bonds, the economy’s reserves bymillion, and the money supply will bymillion.

Now suppose the Fed lowers the reserve requirement to 15 percent, but banks choose to hold another 5 percent of deposits as excess reserves.

True or False: The money multiplier will decrease.

True

False

True or False: As a result, the overall change in the money supply will decrease.

True

False
Answer #1

Reserve requirement = 20%=0.2

Money multiplier =1/r=1/0.2=5

when Fed sells $3million then money supply will decrease by $15million

when reserve requirement=15% then money multiplier=1/r=1/0.15=6.667 and overall money will further decrease.

1st statement is false because money Multiplier increases.

2nd statement is true because money supply will decrease supply by larger amount